# TFusionOcc: Student’s t-Distribution Based Object-Centric Multi-Sensor Fusion Framework for 3D Occupancy Prediction

### v1

@@ III-G5 Student-T-Distribution Based SuperQuadric With Inverse-Warp Primitive
So far, thanks to the T-distribution kernel enhancement, the T-Superquadric primitive is robust to outliers. Nevertheless, the T-Superquadric primitive has a very limited shape variety. It can only exhibit convex, symmetric shapes and cannot capture non-convex, asymmetric shapes, which are common in the daily 3D driving scene. To solve this limitation, we make the T-Superquadric primitive deformable through inverse-warp enhancement. Specifically, we choose 24 basis field functions $\left\{B_{i}(u,v,w)\right\}_{i=1}^{24}$, where $u,v,w$ refers to three inputs, respectively, to deform the local coordinates $(X_{S},Y_{S},Z_{S})$ as follows:
$u=\frac{X_{S}}{S_{x}},v=\frac{Y_{S}}{S_{y}},w=\frac{Z_{S}}{S_{z}}$ (14)
$\begin{bmatrix}\tilde{X}_{S}\\ \tilde{Y}_{S}\\ \tilde{Z}_{S}\\ \end{bmatrix}=\begin{bmatrix}X_{S}\\ Y_{S}\\ Z_{S}\\ \end{bmatrix}-\sum_{i=1}^{24}\omega_{i}\cdot B_{i}\left(u,v,w\right)$ (15)
@@
where $\omega_{i}$ is the learnable weighting scaler valued between -1 and 1. Then, the occupancy probability of the 3D point associated with each T-Superquadric-Inverse-Warp primitive can be calculated as follows:
$f\left(\tilde{X}_{S},\tilde{Y}_{S},\tilde{Z}_{S}\right)=\left(\left(\frac{\tilde{X}_{S}}{S_{x}}\right)^{\frac{2}{\epsilon_{2}}}+\left(\frac{\tilde{Y}_{S}}{S_{y}}\right)^{\frac{2}{\epsilon_{2}}}\right)^{\frac{\epsilon_{2}}{\epsilon_{1}}}+\left(\frac{\tilde{Z}_{S}}{S_{z}}\right)^{\frac{2}{\epsilon_{1}}}$ (16)
$p\left(x|P^{T}_{SQIWi}\right)=\left(1+\frac{1}{\nu}\cdot f\left(\tilde{X}_{S},\tilde{Y}_{S},\tilde{Z}_{S}\right)\right)^{-\frac{\nu+3}{2}}$ (17)
@@
TABLE I: The 24 basis field functions make the T-Superquadric deformable. Those functions enable full shears, twists, bends, quadratics, radial bulge and smooth corners.
III-H Transformer
The general inner structure of the proposed transformer is shown in Figure 2. It takes a set of T-primitives, multi-scale lifted depth-aware visual features, and a semantic-aware lidar feature as input, and proceeds through a fused depth-map-guided 3D deformable attention module, a multi-gate cross-attention fusion module, two MLP-based feed-forward modules, a 3D sparse convolution-based self-attention module, and a T-primitive properties refine module iteratively $N$ times, resulting in the last refined T-primitives. The following sections will introduce the key implementation details of each module.
III-H1 Fused Depth Map based 3D Deformable Attention
@@
Figure 7: 3D Deformable Attention For the same visual feature at location $(u_{i},v_{j})$, several different depth values ranging between $[0,1.0]$ along the depth-axis are used to weight the same pixel feature, resulting in the depth encoded visual feature to mitigate the projection ambiguity problem stemming from the projection process.
@@
Taking multi-scale lifted depth-aware visual features and a set of T-primitive as input, we use a 3D deformable attention operator, namely DFA3D [11], which performs feature aggregation on lifted visual features to mitigate the 3D to 2D projection ambiguity problem, to aggregate visual features for each T-primitive. As shown in Figure 2 camera branch and Figure 7, the multi-scale lifted depth-aware visual features $Vol^{Cam}_{depth}$ are obtained through outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$, this operation gives an accurate depth encoding for each visual feature in the feature map. For each T-primitive feature sampling $p^{T}_{i}$, we first use the mean location $m_{i}$ of each $p^{T}_{i}$ as the center, convert the center location from the cylindrical coordinate to the cartesian coordinate, and feed its associate query feature $q_{i}$ to an MLP layer, resulting in 13 offsets $\Delta m_{j}$ with respect to the center to form a set of 3D sampling points $m^{samp}=\left\{m^{samp}_{i}=m_{i}+\Delta m_{j}\mid j=1,...,13\right\}$. Then, we leverage extrinsic and intrinsic matrices to project each 3D sampling point $m^{samp}_{i}$ onto the lifted depth-aware visual features, resulting in the location of the sampling of the pixel frame $\bar{m}^{samp}_{i}=\left(u_{i},v_{i},d_{i}\right)$. Lastly, a 3D deformable attention operator DFA3D takes $\bar{m}^{samp}_{i}$ as a sampling location to perform cross-attention visual feature aggregation on $Vol^{Cam}_{depth}$, resulting in an updated query feature $q_{i}^{updated}$. The overall sampling process can be formulated as follows:
$q_{i}^{updated}=\sum_{i=1}^{13}DFA3D\left(q_{i},\bar{m}^{samp}_{i},Vol^{Cam}_{depth}\right)$ (18)
@@
After the 3D cross-attention process, $q_{i}^{updated}$ already aggregates sufficient depth-aware visual features and is passed to the MGCAFusion module for feature fusion. Meanwhile, the set of 3D sampling points $m^{samp}$ for each T-primitive is also passed on to the MGCAFusion module for semantic-aware lidar feature aggregation.
III-H2 Multi-Gate Cross-Attention Fusion (MGCAFusion) Module
@@
Figure 8: Multi-Gate Cross Attention Fusion Module. The left part exhibits the multi-modality weighted summation fusion part, and the right part exhibits the multi-modality gated concatenation fusion part. The outputs of the two major parts are further fused to produce the final fused feature.
@@
Given a set of 3D sampling points $m^{samp}$ for each T-primitive, we apply bilinear interpolation on $Vol^{Lidar}_{sem}$ to aggregate a semantic-aware lidar feature for each T-primitive. The detailed implementation of the multi-gate cross attention fusion module (MGCAFusion) is demonstrated in Figure 8. It consists of two central parts, the weighted summation fusion part, as highlighted in the green rectangle on the left of Figure 8 and the gated concatenation fusion part, as highlighted in the pink rectangle on the right of Figure 8. Both parts take the same inputs, but perform different feature fusion strategies.
@@
For the weighted summation fusion part, the semantic-aware lidar feature and the depth-aware visual feature are summed in elements first, then passed through an MLP layer followed by a softmax layer, resulting in two weight values, $W_{Lidar}$ and $W_{Vis}$. The $W_{Lidar}$ and $W_{Vis}$ are multiplied again with the corresponding semantic-aware lidar feature and depth-aware visual feature, respectively, resulting in the weighted lidar feature and the weighted visual feature. Lastly, the weighted lidar feature and the weighted visual feature are summed and passed through an MLP layer, yielding the summation fusion feature.
@@
For the gated concatenation fusion part, the semantic-aware lidar feat and the depth-aware visual feat are passed through an MLP layer, followed by a sigmoid function, respectively, resulting in a visual gate vector $G_{Vis}$ and a lidar gate vector $G_{Lidar}$. $G_{Vis}$ is element-wise multiplied with the semantic-aware lidar feature, followed by a skip connection that results in the gated lidar feature. $G_{Lidar}$ is element-wise multiplied with the depth-aware visual feature, followed by a skip connection that results in the gated visual feature. The gated Lidar and visual features are concatenated along the feature channel dimension, followed by an MLP layer that reduces the fused feature channel dimension, yielding the concatenation fusion feature.
@@
Lastly, the summation and concatenation fusion features are combined via feature-channel concatenation, followed by an MLP layer, yielding the final fused feature.
III-H3 3D Sparse Conv-based Self-Attention Module
Figure 9: Sparse-Convolution-Based Self-Attention Pipeline. The center of T-primitives is treated as a 3D point to form a primitive-based 3D point cloud. We apply a cylindrical partition to the obtained 3D point cloud, followed by a 3D sparse convolution to realise self-attention.
@@
The self-attention module is demonstrated in Figure 9. Each T-primitive center $m_{i}$ is treated as a 3D point, and we can acquire a set of 3D points based on T-primitive. Then, this T-primitive-based sparse 3D point cloud undergoes a coordinate transformation from Cartesian to cylindrical coordinates, followed by voxelization, yielding a sparse voxel volume in cylindrical coordinates. Lastly, the 3D sparse convolution applies to this sparse voxel volume to perform the self-attention operation. Since our T-primitives are initialized in cylindrical coordinates, we perform the self-attention operation in the same coordinates to better preserve the scene’s fine-grained 3D geometry.
III-H4 T-Primitive Property Refinement Module
Since T-primitive queries have aggregated sufficient semantic and 3D geometry information in the previous 3D deformable attention, self-attention, and MGCAFusion modules, we use each query vector to update its associated T-primitive properties. Specifically, for three types of T-primitives, we feed each type of T-primitive query to an MLP layer to obtain the intermediate properties, and we combine the intermediate properties with the initial properties to achieve the refinement of the T-primitive property. The property refinement details for three types of T-primitives are formulated as:
$MLP(Q_{P_{i}})\Longrightarrow(\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{c}_{i})$
$MLP(Q_{SQ_{i}})\Longrightarrow(\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})$
$MLP(Q_{SQIW_{i}})\Longrightarrow(\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{w}_{i},\hat{c}_{i})$
$P^{T}_{New_{i}}=(m+\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{c}_{i})$
$P^{T}_{New_{SQ_{i}}}=(m+\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})$
$P^{T}_{New_{SQIW_{i}}}=(m+\hat{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{w}_{i},\hat{c}_{i})$ (19)
@@ IV-A Implementation Details
The TFusionOcc model takes six surround-view images and 10 lidar sweeps per data sample and leverages ResNet50-DCN [5] as a 2D Backbone, initialized with weights from FCOS3D [28]. The transformer layer is stacked four times iteratively. AdamW optimizer is used, with an initial learning rate of 2e-4 and a weight decay of 0.01. The learning rate is decayed using a multistep scheduler. The predicted occupancy has a resolution of $200\times 200\times 16$ for full-scale evaluation on nuScenes-Surroundocc, nuScenes-Occ3D and nuScenes-C datasets. For data augmentation, we employ photometric distortions and a grid mask to the input images during training. Model training is conducted on two H100 GPUs with 94GB of memory.
@@ IV-B Loss Function
To train TFusionOcc, we take advantage of the Lovász-Softmax loss [1] $L_{lovasz}$ and the binary cross-entropy loss $L_{BCE}$ to optimize the 3D semantic occupancy task. For depth supervision, a binary cross-entropy loss $L^{depth}_{BCE}$ is used to refine the depth distribution feature. The final loss is formulated as follows.
@@ IV-C Dataset
To perform the 3D semantic occupancy prediction task, we use dense labels obtained from SurroundOcc [30] and Occ3D [27], respectively. Since the test set lacks semantic labels, we train our model on the training set and evaluate its performance using the validation set. For annotations from SurroundOcc, we set the range of the X and Y axes to [-50, 50] meters and the Z axis to [-5, 3] meters under lidar coordinates. For annotations from Occ3D, we set the range of the X and Y axes to [-40, 40] meters and the Z axis to [-1, 5.4] meters under the ego coordinates. The input images have a resolution of $1600\times 900$ pixels, while the final output of the semantic occupancy prediction branch is represented with a resolution of $200\times 200\times 16$. The annotations from SurroundOcc contain a total of 17 semantic classes, while the annotations from Occ3D comprise 18 semantic classes. Label 0 refers to free voxels in both annotations.
Additionally, following the methodology proposed in [18], we conduct an in-depth analysis of our model performance in challenging scenarios, specifically in rainy and night conditions. This evaluation is carried out using the annotation file provided by [18].
IV-D Performance Evaluate Metrics
To assess the performance of various SOTA algorithms and compare them with our approach in the 3D semantic occupancy prediction task, we utilize the intersection over union (IoU) to evaluate each semantic class. Moreover, we employ the mean IoU over all semantic classes (mIoU) as a comprehensive evaluation metric:
$IoU=\frac{TP}{TP+FP+FN}$ (21)
and
$mIoU=\frac{1}{Cls}\sum_{i=1}^{Cls}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ (22)
where $TP$, FPFP, and $FN$ represent the counts of true positives, false positives, and false negatives in our predictions, respectively, while $Cls$ denotes the total class number.
IV-E Model Performance Analysis

### v2

@@ III-F5 Student-T-Distribution Based SuperQuadric With Inverse-Warp Primitive
So far, thanks to the T-distribution kernel enhancement, the T-Superquadric primitive is robust to outliers. Nevertheless, due to limited shape variety, it cannot capture non-convex, asymmetric shapes, which are very common in the daily 3D driving scene. To solve this limitation, we make the T-Superquadric primitive deformable through inverse-warp enhancement. Specifically, we choose 24 basis field functions $\left\{B_{k}(u,v,w)\right\}_{k=1}^{24}$ to deform the local coordinates $(X_{S_{i}},Y_{S_{i}},Z_{S_{i}})$ as follows:
$u=\frac{X_{S_{i}}}{S_{x_{i}}},v=\frac{Y_{S_{i}}}{S_{y_{i}}},w=\frac{Z_{S_{i}}}{S_{z_{i}}}$ (13)
$\begin{bmatrix}\tilde{X}_{S_{i}}\\ \tilde{Y}_{S_{i}}\\ \tilde{Z}_{S_{i}}\\ \end{bmatrix}=\begin{bmatrix}X_{S_{i}}\\ Y_{S_{i}}\\ Z_{S_{i}}\\ \end{bmatrix}-\sum_{k=1}^{24}\omega_{i,k}\cdot B_{k}\left(u,v,w\right)$ (14)
@@
Then, the occupancy probability of the 3D point associated with $i$-th T-Superquadric-Inverse-Warp primitive can be calculated as follows:
$f\left(\tilde{X}_{S_{i}},\tilde{Y}_{S_{i}},\tilde{Z}_{S_{i}}\right)=\left(\left(\frac{\tilde{X}_{S_{i}}}{S_{x_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}+\left(\frac{\tilde{Y}_{S_{i}}}{S_{y_{i}}}\right)^{\frac{2}{\epsilon_{2_{i}}}}\right)^{\frac{\epsilon_{2_{i}}}{\epsilon_{1_{i}}}}+\left(\frac{\tilde{Z}_{S_{i}}}{S_{z_{i}}}\right)^{\frac{2}{\epsilon_{1_{i}}}}$ (15)
$p\left(x|P^{T}_{SQIWi}\right)=\left(1+\frac{1}{\nu_{i}}\cdot f\left(\tilde{X}_{S_{i}},\tilde{Y}_{S_{i}},\tilde{Z}_{S_{i}}\right)\right)^{-\frac{\nu_{i}+3}{2}}$ (16)
TABLE I: The 24 basis field functions make the T-Superquadric deformable. Those functions enable full shears, twists, bends, quadratics, radial bulge, and smooth corners.
@@
III-G Transformer
III-G1 Fused Depth Map Guided 3D Deformable Attention
@@
Figure 7: Fused Depth Map Guided 3D Deformable Attention For the same visual feature at location $(u_{i},v_{j})$, several different depth values ranging between $[0.0,1.0]$ along the depth-axis are used to weight the same pixel feature, resulting in the depth encoded visual feature to mitigate the projection ambiguity problem stemming from the projection process.
@@
Taking multi-scale lifted depth-aware visual features and a set of T-primitive as input, we use a 3D deformable attention operator, namely 3DFA [11], which performs feature aggregation on lifted visual features to mitigate the 3D to 2D projection ambiguity problem, to aggregate visual features for each T-primitive. As shown in Fig. 2 camera branch and Fig. 7, the multi-scale lifted depth-aware visual features $Vol^{Cam}_{depth}$ are obtained through outer product operation $F_{depth}^{Cam}\otimes V_{depth}^{fuse}$, this operation gives an accurate depth encoding for each visual feature in the feature map. For each T-primitive feature sampling $p^{T}_{i}$, we first use the mean location $m_{i}$ of each $p^{T}_{i}$ as the center, convert the center location from the cylindrical coordinate to the cartesian coordinate, and feed its associate query feature $q_{i}$ to an MLP layer, resulting in 13 offsets $\Delta m_{j}$ with respect to the center to form a set of 3D sampling points $m^{samp}=\left\{m^{samp}_{i}=m_{i}+\Delta m_{j}\mid j=1,...,13\right\}$. Then, we leverage extrinsic and intrinsic matrices to project each 3D sampling point $m^{samp}_{i}$ onto the lifted depth-aware visual features, resulting in the location of the sampling of the pixel frame $\bar{m}^{samp}_{i}=\left(u_{i},v_{i},d_{i}\right)$. Lastly, a 3D deformable attention operator 3DFA takes $\bar{m}^{samp}_{i}$ as a sampling location to perform cross-attention visual feature aggregation on $Vol^{Cam}_{depth}$, resulting in an updated query feature $q_{i}^{updated}$. The overall sampling process can be formulated as follows.
$q_{i}^{updated}=\sum_{i=1}^{13}3DFA\left(q_{i},\bar{m}^{samp}_{i},Vol^{Cam}_{depth}\right)$ (17)
@@
After the 3D cross-attention process, $q_{i}^{updated}$ already aggregates sufficient depth-aware visual features and is passed to the GCWSFusion module for feature fusion. Meanwhile, the set of 3D sampling points $m^{samp}$ for each T-primitive is also passed on to the GCWSFusion module for semantic-aware LiDAR feature aggregation.
III-G2 Gate-Concatenation and Weight-Summation Fusion (GCWSFusion) Module
@@
Figure 8: Gate-Concatenation and Weight-Summation Fusion Module. The left part exhibits the multi-modality weighted summation fusion part, and the right part exhibits the multi-modality gated concatenation fusion part. The outputs of the two major parts are further fused to produce the final fused feature for each T-Primitive.
@@
Given a set of 3D sampling points $m^{samp}$ for each T-primitive, we apply bilinear interpolation on $Vol^{Lidar}_{sem}$ to aggregate a semantic-aware LiDAR feature for each T-primitive. Then, both the semantic-aware LiDAR feature and the depth-aware visual feature of each T-primitive are fed into the GCWSFusion module for late-stage feature fusion. The detailed implementation of GCWSFusion module is demonstrated in Fig. 8. It consists of two central parts, the weighted summation fusion part, as highlighted in the green rectangle on the left of Fig. 8, and the gated concatenation fusion part, as highlighted in the pink rectangle on the right of Fig. 8. Both parts take the same inputs, but perform different feature fusion strategies.
@@
For the weighted summation fusion part, the semantic-aware LiDAR feature and the depth-aware visual feature are summed in elements first, then passed through several linear layers followed by a softmax layer, resulting in two weight values, $W_{Lidar}$ and $W_{Vis}$. The $W_{Lidar}$ and $W_{Vis}$ are multiplied again with the corresponding semantic-aware LiDAR feature and depth-aware visual feature, respectively, resulting in the weighted LiDAR feature and the weighted visual feature. Lastly, the weighted LiDAR feature and the weighted visual feature are summed and passed through an extra set of linear layers, yielding the summation fusion feature. This design adaptively weights the dominance information across two sensors.
@@
For the gated concatenation fusion part, the semantic-aware LiDAR feat and the depth-aware visual feat are passed through several linear layers, followed by a sigmoid function, respectively, resulting in a visual gate vector $G_{Vis}$ and a LiDAR gate vector $G_{Lidar}$. $G_{Vis}$ is element-wise multiplied with the semantic-aware LiDAR feature, followed by a skip connection that results in the gated LiDAR feature. $G_{Lidar}$ is element-wise multiplied with the depth-aware visual feature, followed by a skip connection that results in the gated visual feature. The gated LiDAR and visual features are concatenated along the feature channel dimension, followed by another set of linear layers that reduces the fused feature channel dimension, yielding the concatenation fusion feature. This design can dynamically amplify or suppress feature information within the two different modalities, thereby achieving denoising and information enhancement during the feature fusion process.
@@
Lastly, the summation and concatenation fusion features are combined via feature-channel concatenation, followed by extra linear layers, yielding the final fused feature.
III-G3 3D Sparse Conv-based Self-Attention Module
@@
Each T-primitive center $m_{i}$ is treated as a 3D point, and we can acquire a set of 3D points based on a set of T-primitives. Then, this T-primitive-based sparse 3D point cloud undergoes a coordinate transformation from Cartesian to Cylindrical coordinates, followed by voxelization, yielding a sparse voxel volume in cylindrical coordinates. Lastly, the 3D sparse convolution applies to this sparse voxel volume to perform the self-attention operation. Since our T-primitives are initialized in cylindrical coordinates, we perform the self-attention operation in the same coordinates to better preserve the scene’s fine-grained 3D geometry.
III-G4 T-Primitive Property Refinement
As presented in (18), the physical attributes of the three proposed T-Primitives are refined by processing each associated query vector through an MLP to produce updated attributes, which are subsequently added to the original attributes for refinement.
$(\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{c}_{i})=MLP(Q_{P_{i}})$
$(\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})=MLP(Q_{SQ_{i}})$
$(\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\left\{\hat{\omega_{i,k}}\right\}_{k=1}^{24},\hat{c}_{i})=MLP(Q_{SQIW_{i}})$
$\hat{P^{T}_{i}}=(m_{i}+\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{c}_{i})$
$\hat{P^{T}_{SQ_{i}}}=(m_{i}+\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\hat{c}_{i})$
$\hat{P^{T}_{SQIW_{i}}}=(m_{i}+\Delta{m}_{i},\hat{s}_{i},\hat{r}_{i},\hat{\alpha}_{i},\hat{\epsilon}_{1_{i}},\hat{\epsilon}_{2_{i}},\left\{\hat{\omega_{i,k}}\right\}_{k=1}^{24},\hat{c}_{i})$ (18)
@@ IV-A Implementation Details
The TFusionOcc model takes six surround-view images and 10 LiDAR sweeps per data sample and leverages ResNet50-DCN [5] as a 2D Backbone, initialized with weights from FCOS3D [29]. The transformer layer is stacked four times iteratively. AdamW optimizer is used, with an initial learning rate of 2e-4 and a weight decay of 0.01. The learning rate is decayed using a multistep scheduler. The predicted occupancy has a resolution of $200\times 200\times 16$ for full-scale evaluation on SurroundOcc-nuScenes, Occ3D-nuScenes and nuScenes-C datasets. For data augmentation, we employ photometric distortions and a grid mask to the input images during training. Model training is conducted on two H100 GPUs with 94GB of memory.
@@ IV-B Loss Function
To train TFusionOcc, we take advantage of the Lovász-Softmax loss [1] $L_{lovasz}$ and the binary cross-entropy loss $L_{BCE}$ to optimize the 3D semantic occupancy task. For depth supervision, a binary cross-entropy loss $L^{depth}_{BCE}$ is used to refine the depth distribution feature. The final loss is formulated as:
@@ IV-C Dataset
To perform the 3D semantic occupancy prediction task, we use dense labels obtained from SurroundOcc [31] and Occ3D [28], respectively. Since the test set lacks semantic labels, we train our model on the training set and evaluate its performance using the validation set. For SurroundOcc annotations, we set the range of the X and Y axes to [-50, 50] meters and the Z axis to [-5, 3] meters under LiDAR coordinates. For annotations from Occ3D, we set the range of the X and Y axes to [-40, 40] meters and the Z axis to [-1, 5.4] meters under the ego coordinates. The input images have a resolution of $1600\times 900$ pixels, while the final output of the semantic occupancy prediction branch is represented with a resolution of $200\times 200\times 16$. The annotations from SurroundOcc contain a total of 17 semantic classes, while the annotations from Occ3D comprise 18 semantic classes. Label 0 refers to free voxels in both annotations.
IV-D Model Performance Analysis